#4. Jason Matthews purchased 100 shares of the Hodge & Mattox Energy Fund. Each share cost $14.15. Fifteen months later, he decided to sell his shares when the share value reached $16.80

a) What was the amount of his total investment?
b) What was the total amount Jason received when he sold his shares to the Hodge fund?
c) How much profit did he make on his investment?
Answer #1

a)Amount of total investment = Number of shares*Cost per share

= 100*14.15

= $1,415

b)Total amount received on sale = Number of shares*Sale Price per share

= 100*16.80

= $1,680

c)Profit = 1,680-1,415

= $265